# Canonical bases of tensor products of integrable highest weight modules arising from framed constructions

### v2

@@ Abstract.
Given a quantum group, we prove that the canonical bases of the tensor products of its integrable highest weight modules can be obtained from the canonical bases of the simple integrable highest modules of a bigger quantum group. As a result, based on the positivity of the canonical bases of the simple integrable highest modules due to Lusztig, we prove that the canonical bases of the tensor products have the positivity.
@@ 1. Introduction
For any symmetric Cartan datum $(I,\cdot)$ and $Y$-regular root datum $(Y,X,\langle,\rangle,...)$ of type $(I,\cdot)$, let $\mathbf{f}$ be the corresponding algebra and $\mathbf{U}$ be the corresponding quantized enveloping algebra defined in [13, Chapter 1, 3]. For any dominant weight $\lambda$, let $\Lambda_{\lambda}$ be the simple integrable highest weight module of $\mathbf{U}$ with the highest weight $\lambda$, see section 2 for details.
@@
In [10, 11], Lusztig categorified the algebra $\mathbf{f}$ by perverse sheaves on the varieties of quiver representations, and constructed the canonical basis $\mathbf{B}$ of $\mathbf{f}$ with some very remarkable properties, such as the integrality, the positivity for the structure constants with respect to the multiplication and the comultiplication of $\mathbf{f}$, and the compatibility with various natural filtrations. In particular, the canonical basis $\mathbf{B}$ naturally gives rise to the canonical basis $\mathbf{B}(\Lambda_{\lambda})$ of $\Lambda_{\lambda}$, see section 2 for details. Lusztig proved that the canonical basis $\mathbf{B}(\Lambda_{\lambda})$ has the positivity for the structure constants with respect to the actions of Chevalley generators $F_{i},E_{i}\in\mathbf{U}$, see [13, Theorem 22.1.7]. We remark that this proof relies on the positivity of $\mathbf{B}$. In [6, 7], Kashiwara constructed the canonical bases $\mathbf{B}$ of $\mathbf{f}$ and $\mathbf{B}(\Lambda_{\lambda})$ of $\Lambda_{\lambda}$ by his crystal basis theory. However, this elegant construction can not provide the positivity. The proof for the positivity of the canonical basis $\mathbf{B}(\Lambda_{\lambda})$ given by the categorification of the module $\Lambda_{\lambda}$ is expected. In [5], Kang-Kashiwara categorified $\Lambda_{\lambda}$ by finitely generated graded projective modules of cyclotomic Khovanov-Lauda-Rouquier algebra. In [18], Zheng categorified $\Lambda_{\lambda}$ by perverse sheaves on the stacks of representations of the framed quiver, and proved the positivity of $\mathbf{B}(\Lambda_{\lambda})$. In [4], based on Zheng’s work, Fang-Lan-Xiao gave a categorification of $\Lambda_{\lambda}$ by Lusztig sheaves for the framed quiver, and proved the positivity of $\mathbf{B}(\Lambda_{\lambda})$. We remark that the framed quiver construction is inspired by Nakajima’s work [14].
@@
In [12], by using of the $-\mathcal{R}$-matrix, Lusztig constructed the canonical basis of the tensor product of a simple integrable lowest weight module followed by a simple integrable highest weight module. In [13, Chapter 27], he generalized this method to construct the canonical basis of the tensor product of based modules when $(I,\cdot)$ is of finite type. In [2], Bao-Wang generalized Lusztig’s method to construct the canonical basis of the tensor product of several integrable lowest weight modules followed by integrable highest weight modules for any $(I,\cdot)$.
@@
It is natural to consider the question whether the canonical basis of the tensor product has the positivity for the structure constants with respect to the actions of Chevalley generators $F_{i},E_{i}\in\mathbf{U}$. There are positive answers for the tensor product of several highest weight modules when $(I,\cdot)$ is symmetric. In [17], Webster categorified the tensor product building on works of [8, 15, 16], and proved the positivity of the canonical basis. In [18], Zheng categorified the tensor product by perverse sheaves on the stacks of representations of the framed quiver. In [3], Fan-Lan categorified the tensor product by Lusztig sheaves for the $N$-framed quiver, and proved the positivity of the canonical basis.
@@
Step 1. For the Cartan datum $(I,\cdot)$ and the root datum $(Y,X,\langle,\rangle,...)$, we enlarge them to be the framed Cartan datum $(\tilde{I},\cdot)$ and the framed root datum $(\tilde{Y},\tilde{X},\langle,\rangle,...)$, see section 3.1 for details. Then we have the corresponding algebra $\tilde{\mathbf{f}}$ and the corresponding quantized enveloping algebra $\tilde{\mathbf{U}}$ such that $\mathbf{f}$ is a subalgebra of $\tilde{\mathbf{f}}$, and $\mathbf{U}$ is a subalgebra of $\tilde{\mathbf{U}}$.
Step 2. For any dominant weights $\xi,\lambda$ with respect to $(Y,X,\langle,\rangle,...)$, we define a dominant weight $\xi\odot\lambda$ with respect to $(\tilde{Y},\tilde{X},\langle,\rangle,...)$, see section 3.2 for details. Then we have the simple integrable highest weight module $\tilde{\Lambda}_{\xi\odot\lambda}$ of $\tilde{\mathbf{U}}$ with the canonical basis $\tilde{\mathbf{B}}(\tilde{\Lambda}_{\xi\odot\lambda})$.
@@
Step 4. By using of the comultiplication of $\tilde{\mathbf{f}}$, we construct a $\mathbf{U}$-module isomorphism $\varphi:\Lambda_{\xi,\lambda}\rightarrow\Lambda_{\xi}\otimes\Lambda_{\lambda}$, see Theorem 3.8. Moreover, we prove that the image of the basis $\tilde{\mathbf{B}}(\Lambda_{\xi,\lambda})$ under $\varphi$ coincides with the canonical basis of the tensor product $\Lambda_{\xi}\otimes\Lambda_{\lambda}$, see Theorem 4.7.
@@
We remark that the $\mathbf{U}$-module $\Lambda_{\xi,\lambda}$ has been constructed by Li in [9, Section 3.2, 3.3]. Moreover, he constructed a $\mathbf{U}$-module isomorphism $\Lambda_{\xi}\otimes\Lambda_{\lambda}\rightarrow\Lambda_{\xi,\lambda}$. This isomorphism is rather different from the isomorphism $\varphi:\Lambda_{\xi,\lambda}\rightarrow\Lambda_{\xi}\otimes\Lambda_{\lambda}$ in Step 4. By using of the construction of $\varphi$, we can prove the positivity of the transition matrix between the bases $\mathbf{B}(\Lambda_{\xi}\otimes\Lambda_{\lambda})$ and $\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$, see Corollary 4.8. This positivity result and the positivity of the canonical basis $\mathbf{B}(\Lambda_{\xi}\otimes\Lambda_{\lambda})$ are not contained in [9].
@@ Acknowledgments
J. Fang is supported by the New Cornerstone Science Foundation through the New Cornerstone Investigator Program awarded to Professor Xuhua He, and the National Natural Science Foundation of China (No. 12471030). Y. Lan is supported by the National Natural Science Foundation of China (No. 1288201, 12471030). The authors would like to thank Professor Jie Xiao for his helpful discussions and suggestions. The authors would like to thank Professor Huanchen Bao for pointing out that our idea in [3] could be combined with [13, Theorem 22.1.7] due to Lusztig to give an algebraic proof for the positivity which is independent of the geometric methods in [3].
@@ 2. Preliminaries
Throughout this paper, we fix a symmetric Cartan datum $(I,\cdot)$ and a $Y$-regular root datum $(Y,X,\langle,\rangle,...)$ of type $(I,\cdot)$. More precisely, we fix a finite set $I$; a symmetric bilinear form $\mathbb{Z}[I]\times\mathbb{Z}[I]\rightarrow\mathbb{Z},(\nu,\nu^{\prime})\mapsto\nu\cdot\nu^{\prime}$; two finitely generated free abelian groups $Y$ and $X$; a perfect bilinear pairing $\langle,\rangle:Y\times X\rightarrow\mathbb{Z}$; two embeddings $I\hookrightarrow Y,i\mapsto i$ and $I\hookrightarrow X,i\mapsto i^{*}$, such that
@@
The embeddings $I\hookrightarrow Y$ and $I\hookrightarrow X$ induce homomorphisms $\mathbb{Z}[I]\rightarrow Y$ and $\mathbb{Z}[I]\rightarrow X$, we shall often denote the images of $\nu\in\mathbb{Z}[I]$ under either of these homomorphisms by the same notation $\nu$.
@@
$[n]=\frac{v^{n}-v^{-n}}{v-v^{-1}},[m]!=\prod^{m}_{k=1}[k],\begin{bmatrix}n\\ m\end{bmatrix}=\frac{\prod_{k=0}^{m-1}(v^{n-k}-v^{-n+k})}{\prod_{k=1}^{m}(v^{k}-v^{-k})}.$
@@ 2.1. The algebra $\mathbf{f}$
The bar-involution of $\mathbf{f}$ is the unique $\mathbb{Q}$-algebra involution ¯$\bar{}:\mathbf{f}\rightarrow\mathbf{f}$ such that
@@
There is a unique non-degenerate symmetric bilinear inner product $(,):\mathbf{f}\times\mathbf{f}\rightarrow\mathbb{Q}(v)$ such that
@@
(b) $(x_{1}x_{2},x)=(x_{1}\otimes x_{2},r(x))$ for any $x_{1},x_{2},x\in\mathbf{f}$, where the inner product on $\mathbf{f}\otimes\mathbf{f}$ is given by $(x_{1}\otimes x_{2},x^{\prime}_{1}\otimes x^{\prime}_{2})=(x_{1},x^{\prime}_{1})(x_{2},x^{\prime}_{2})$ for any $x_{1},x_{2},x^{\prime}_{1},x^{\prime}_{2}\in\mathbf{f}$.
@@ 2.2. The algebra $\mathbf{U}$
For the root datum $(Y,X,\langle,\rangle,...)$, the quantized enveloping algebra $\mathbf{U}$ is the $\mathbb{Q}(v)$-algebra with $1$ generated by $\{E_{i},F_{i},K_{\mu}\mid i\in I,\mu\in Y\}$ subject to the following relations
@@
The bar-involution of $\mathbf{U}$ is the unique $\mathbb{Q}$-algebra involution ¯$\bar{}:\mathbf{U}\rightarrow\mathbf{U}$ such that
@@ 2.3. The module $\Lambda_{\lambda}$
A $\mathbf{U}$-module $M$ is a weight module, if it has a decomposition $M=\bigoplus_{\lambda\in X}M^{\lambda}$ such that $K_{\mu}m=v^{\langle\mu,\lambda\rangle}m$ for any $\mu\in Y,\lambda\in X$ and $m\in M^{\lambda}$. If $M^{\lambda}\not=0$, then $M^{\lambda}\not=0$ is called a weight subspace of $M$. Throughout this paper, we only consider weight modules.
@@
A symmetric bilinear inner product $(,):M\times M\rightarrow\mathbb{Q}(v)$ is called admissible, if
@@
For any dominant weight $\lambda\in X$, that is, $\langle i,\lambda\rangle\geqslant 0$ for any $i\in I$, let $\sum_{i\in I}\mathbf{f}\theta_{i}^{\langle i,\lambda\rangle+1}$ be the left ideal of $\mathbf{f}$ generated by $\{\theta_{i}^{\langle i,\lambda\rangle+1}\mid i\in I\}$. This left ideal is a submodule of $M_{\lambda}$, see [13, Proposition 3.5.6(a)]. The simple integrable highest weight module $\Lambda_{\lambda}$ is the quotient
@@
The bar-involution of $\Lambda_{\lambda}$ is the unique $\mathbb{Q}$-linear involution ¯$\bar{}:\Lambda_{\lambda}\rightarrow\Lambda_{\lambda}$ such that
@@
There is a unique admissible symmetric bilinear inner product $(,)_{\lambda}:\Lambda_{\lambda}\times\Lambda_{\lambda}\rightarrow\mathbb{Q}(v)$ such that $(\eta_{\lambda},\eta_{\lambda})_{\lambda}=1$.
@@ 2.4. The tensor product $\Lambda_{\xi}\otimes\Lambda_{\lambda}$
For any two dominants weights $\xi,\lambda\in X$, the tensor product $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ is naturally a $(\mathbf{U}\otimes\mathbf{U})$-module such that
@@
The bar-involution of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ is the $\mathbb{Q}$-linear involution ¯$\bar{}:\Lambda_{\xi}\otimes\Lambda_{\lambda}\rightarrow\Lambda_{\xi}\otimes\Lambda_{\lambda}$ such that
@@
We define a symmetric bilinear inner product $(,)_{\xi,\lambda}:(\Lambda_{\xi}\otimes\Lambda_{\lambda})\times(\Lambda_{\xi}\otimes\Lambda_{\lambda})\rightarrow\mathbb{Q}(v)$ by
@@
By [13, lemma 17.1.3(b)], the inner product $(,)_{\xi,\lambda}$ is admissible.
@@ 2.5. Canonical basis
A basis $B$ of a $\mathbb{Q}(v)$-vector space $V$ is called almost orthogonal with respect to a symmetric bilinear inner product $(,):V\times V\rightarrow\mathbb{Q}(v)$, if
$(b,b^{\prime})\in\delta_{b,b^{\prime}}+v^{-1}\mathbb{Z}[[v^{-1}]]\cap\mathbb{Q}(v)\ \textrm{for any $b,b^{\prime}\in B$}.$
@@
Let $B$ be an almost orthogonal basis of a $\mathbb{Q}(v)$-vector space $V$ with respect to a symmetric bilinear inner product $(,):V\times V\rightarrow\mathbb{Q}(v)$, ${}_{\mathcal{A}}V$ be the $\mathcal{A}$-submodule of $V$ generated by $B$ and $L(V)=\{x\in V\mid(x,x)\in\mathbf{A}\}$.

### v3

@@ Abstract.
Given a quantum group, we prove that the canonical bases of the tensor products of its integrable highest weight modules can be obtained from the canonical bases of the integrable highest weight modules of a bigger quantum group. As a result, based on the positivity of the canonical bases of the integrable highest weight modules due to Lusztig, we prove that the canonical bases of the tensor products have the positivity.
@@ 1. Introduction
[…]
@@
In [10, 11], Lusztig categorified the algebra $\mathbf{f}$ by perverse sheaves on the varieties of quiver representations, and constructed the canonical basis $\mathbf{B}$ of $\mathbf{f}$ with some very remarkable properties, such as the integrality, the positivity for the structure constants with respect to the multiplication and the comultiplication of $\mathbf{f}$, and the compatibility with various natural filtrations. In particular, the canonical basis $\mathbf{B}$ naturally gives rise to the canonical basis $\mathbf{B}(\Lambda_{\lambda})$ of $\Lambda_{\lambda}$, see section 2 for details. Lusztig proved that the canonical basis $\mathbf{B}(\Lambda_{\lambda})$ has the positivity for the structure constants with respect to the actions of Chevalley generators $F_{i},E_{i}\in\mathbf{U}$, see [13, Theorem 22.1.7]. We remark that this proof relies on the positivity of $\mathbf{B}$. In [6, 7], Kashiwara constructed the canonical bases $\mathbf{B}$ of $\mathbf{f}$ and $\mathbf{B}(\Lambda_{\lambda})$ of $\Lambda_{\lambda}$ by his crystal basis theory. However, this elegant construction does not provide the positivity. The proof for the positivity of the canonical basis $\mathbf{B}(\Lambda_{\lambda})$ given by the categorification of the module $\Lambda_{\lambda}$ is expected. In [5], Kang-Kashiwara categorified $\Lambda_{\lambda}$ by finitely generated graded projective modules of cyclotomic Khovanov-Lauda-Rouquier algebra. In [18], Zheng categorified $\Lambda_{\lambda}$ by perverse sheaves on the stacks of representations of the framed quiver, and proved the positivity of $\mathbf{B}(\Lambda_{\lambda})$. In [4], based on Zheng’s work, Fang-Lan-Xiao gave a categorification of $\Lambda_{\lambda}$ by Lusztig sheaves for the framed quiver, and proved the positivity of $\mathbf{B}(\Lambda_{\lambda})$. We remark that the framed quiver construction is inspired by Nakajima’s work [14].
@@
In [12], by using the quasi-$\mathcal{R}$-matrix, Lusztig constructed the canonical basis of the tensor product of a simple integrable lowest weight module followed by an integrable highest weight module. In [13, Chapter 27], he generalized this method to construct the canonical basis of the tensor product of based modules when $(I,\cdot)$ is of finite type. In [2], Bao-Wang generalized Lusztig’s method to construct the canonical basis of the tensor product of several integrable lowest weight modules followed by integrable highest weight modules for any $(I,\cdot)$.
@@
It is natural to consider the question whether the canonical basis of the tensor product has the positivity for the structure constants with respect to the actions of Chevalley generators $F_{i},E_{i}\in\mathbf{U}$. There are positive answers for the tensor product of several highest weight modules when $(I,\cdot)$ is symmetric. In [17], Webster categorified the tensor product building on works of [8, 15, 16], and proved the positivity of the canonical basis. In [18], Zheng categorified the tensor product by perverse sheaves on the stacks of representations of the framed quiver. In [3], Fang-Lan categorified the tensor product by Lusztig sheaves for the $N$-framed quiver, and proved the positivity of the canonical basis.
@@
[…]
@@
Step 4. By using the comultiplication of $\tilde{\mathbf{f}}$, we construct a $\mathbf{U}$-module isomorphism $\varphi:\Lambda_{\xi,\lambda}\rightarrow\Lambda_{\xi}\otimes\Lambda_{\lambda}$, see Theorem 3.8. Moreover, we prove that the image of the basis $\tilde{\mathbf{B}}(\Lambda_{\xi,\lambda})$ under $\varphi$ coincides with the canonical basis of the tensor product $\Lambda_{\xi}\otimes\Lambda_{\lambda}$, see Theorem 4.7.
@@
We remark that the $\mathbf{U}$-module $\Lambda_{\xi,\lambda}$ has been constructed by Li in [9, Section 3.2, 3.3]. Moreover, he constructed a $\mathbf{U}$-module isomorphism $\Lambda_{\xi}\otimes\Lambda_{\lambda}\rightarrow\Lambda_{\xi,\lambda}$. This isomorphism is rather different from the isomorphism $\varphi:\Lambda_{\xi,\lambda}\rightarrow\Lambda_{\xi}\otimes\Lambda_{\lambda}$ in Step 4. By using the construction of $\varphi$, we can prove the positivity of the transition matrix between the bases $\mathbf{B}(\Lambda_{\xi}\otimes\Lambda_{\lambda})$ and $\mathbf{B}(\Lambda_{\xi})\otimes\mathbf{B}(\Lambda_{\lambda})$, see Corollary 4.8. This positivity result and the positivity of the canonical basis $\mathbf{B}(\Lambda_{\xi}\otimes\Lambda_{\lambda})$ are not contained in [9].
@@ Acknowledgments
This work was supported by the New Cornerstone Science Foundation through the New Cornerstone Investigator Program awarded to Professor Xuhua He; and the National Natural Science Foundation of China [Grant numbers 1288201,12471030]. The authors would like to thank Professor Jie Xiao for his helpful discussions and suggestions. The authors would like to thank Professor Huanchen Bao for pointing out that our idea in [3] could be combined with [13, Theorem 22.1.7] due to Lusztig to give an algebraic proof for the positivity which is independent of the geometric methods in [3]. The authors are grateful to the referees for helpful suggestions and corrections.
@@ 2. Preliminaries
[…]
@@
The embeddings $I\hookrightarrow Y$ and $I\hookrightarrow X$ induce homomorphisms $\mathbb{Z}[I]\rightarrow Y$ and $\mathbb{Z}[I]\rightarrow X$, and we shall often denote the images of $\nu\in\mathbb{Z}[I]$ under either of these homomorphisms by the same notation $\nu$.
@@
$n]=\frac{v^{n}-v^{-n}}{v-v^{-1}},[m]!=\prod^{m}_{k=1}[k],\begin{bmatrix}n\\ m\end{bmatrix}=\frac{\prod_{k=0}^{m-1}(v^{n-k}-v^{-n+k})}{\prod_{k=1}^{m}(v^{k}-v^{-k})}.$
@@ 2.1. The algebra $\mathbf{f}$
The bar-involution of $\mathbf{f}$ is the unique $\mathbb{Q}$-algebra involution $\bar{}:\mathbf{f}\rightarrow\mathbf{f}$ such that
@@
There is a unique non-degenerate symmetric bilinear form $(,):\mathbf{f}\times\mathbf{f}\rightarrow\mathbb{Q}(v)$ such that
@@
(b) $(x_{1}x_{2},x)=(x_{1}\otimes x_{2},r(x))$ for any $x_{1},x_{2},x\in\mathbf{f}$, where the bilinear form on $\mathbf{f}\otimes\mathbf{f}$ is given by $(x_{1}\otimes x_{2},x^{\prime}_{1}\otimes x^{\prime}_{2})=(x_{1},x^{\prime}_{1})(x_{2},x^{\prime}_{2})$ for any $x_{1},x_{2},x^{\prime}_{1},x^{\prime}_{2}\in\mathbf{f}$.
@@ 2.2. The algebra $\mathbf{U}$
[…]
@@
The bar-involution of $\mathbf{U}$ is the unique $\mathbb{Q}$-algebra involution $\bar{}:\mathbf{U}\rightarrow\mathbf{U}$ such that
@@ 2.3. The module $\Lambda_{\lambda}$
A $\mathbf{U}$-module $M$ is a weight module, if it has a decomposition $M=\bigoplus_{\lambda\in X}M^{\lambda}$ such that $K_{\mu}m=v^{\langle\mu,\lambda\rangle}m$ for any $\mu\in Y,\lambda\in X$ and $m\in M^{\lambda}$. If $M^{\lambda}\not=0$, then $M^{\lambda}$ is called a weight subspace of $M$. Throughout this paper, we only consider weight modules.
@@
A symmetric bilinear form $(,):M\times M\rightarrow\mathbb{Q}(v)$ is called admissible, if
@@
For any dominant weight $\lambda\in X$, that is, $\langle i,\lambda\rangle\geqslant 0$ for any $i\in I$, let $\sum_{i\in I}\mathbf{f}\theta_{i}^{\langle i,\lambda\rangle+1}$ be the left ideal of $\mathbf{f}$ generated by $\{\theta_{i}^{\langle i,\lambda\rangle+1}\mid i\in I\}$. This left ideal is a submodule of $M_{\lambda}$, see [13, Proposition 3.5.6(a)]. The integrable highest weight module $\Lambda_{\lambda}$ is the quotient
@@
The bar-involution of $\Lambda_{\lambda}$ is the unique $\mathbb{Q}$-linear involution $\bar{}:\Lambda_{\lambda}\rightarrow\Lambda_{\lambda}$ such that
@@
There is a unique admissible symmetric bilinear form $(,)_{\lambda}:\Lambda_{\lambda}\times\Lambda_{\lambda}\rightarrow\mathbb{Q}(v)$ such that $(\eta_{\lambda},\eta_{\lambda})_{\lambda}=1$.
@@ 2.4. The tensor product $\Lambda_{\xi}\otimes\Lambda_{\lambda}$
For any two dominant weights $\xi,\lambda\in X$, the tensor product $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ is naturally a $(\mathbf{U}\otimes\mathbf{U})$-module such that
@@
The bar-involution of $\Lambda_{\xi}\otimes\Lambda_{\lambda}$ is the $\mathbb{Q}$-linear involution $\bar{}:\Lambda_{\xi}\otimes\Lambda_{\lambda}\rightarrow\Lambda_{\xi}\otimes\Lambda_{\lambda}$ such that
@@
We define a symmetric bilinear form $(,)_{\xi,\lambda}:(\Lambda_{\xi}\otimes\Lambda_{\lambda})\times(\Lambda_{\xi}\otimes\Lambda_{\lambda})\rightarrow\mathbb{Q}(v)$ by
@@
By [13, lemma 17.1.3(b)], the bilinear form $(,)_{\xi,\lambda}$ is admissible.
@@ 2.5. Canonical basis
A basis $B$ of a $\mathbb{Q}(v)$-vector space $V$ is called almost orthonormal with respect to a symmetric bilinear form $(,):V\times V\rightarrow\mathbb{Q}(v)$, if
[…]
@@
Let $B$ be an almost orthonormal basis of a $\mathbb{Q}(v)$-vector space $V$ with respect to a symmetric bilinear form $(,):V\times V\rightarrow\mathbb{Q}(v)$, $_{\mathcal{A}}V$ be the $\mathcal{A}$-submodule of $V$ generated by $B$ and $L(V)=\{x\in V\mid(x,x)\in\mathbf{A}\}$.